# DARTH: Declarative Recall Through Early Termination for Approximate Nearest Neighbor Search

### v1

@@ 3.3.2. Integration in IVF
We discuss the implementation of DARTH for the IVF (Douze et al., 2024) index as well, a popular Tree-based ANNS index. IVF performs k-means clustering over the vector collection, generating $nlist$ centroids. Each centroid operates as a bucket, and the collection vectors are placed in the bucket of their nearest centroid. IVF searches through the vectors of the nearest $nprobe$ cluster buckets to search for the nearest neighbors of a query vector.
DARTH can be effectively used for IVF with minimal changes to the input features of Table 1. Specifically, in DARTH for IVF, the $firstNN$ input feature represents the distance of the query to the closest centroid, while the $nstep$ feature represents the number of the cluster bucket we are currently searching. All other input features, as well as the dynamic recall predictor invocations with adaptive intervals, are the same as the HNSW implementation.
@@ 4. Experimental Evaluation
Setup. We conduct our experimental evaluation on a server with Intel® Xeon® E5-2643 v4 CPUs @ 3.40GHz (12 cores, 24 hyperthreads) and 500GB of available main memory. All algorithms are implemented in C/C++, embedded in the FAISS (Douze et al., 2024) library, with SIMD [^2] Single Instruction Multiple Data (SIMD): A parallel computing method where a single instruction operates simultaneously on multiple data points. support for the Euclidean Distance calculations. Our predictor models are implemented using the LightGBM (Ke et al., 2017) library. All implementations are compiled using g++ 11.4.0 on Ubuntu 22.04.4.
@@
Dataset Dimension Base Vectors Description SIFT100M (Jégou et al., 2011) 128 100M Image Descriptors DEEP100M (Babenko and Lempitsky, 2016) 96 100M Deep Image Embeddings T2I100M (Simhadri et al., 2022) 200 100M Image and Text Embeddings GLOVE1M (Pennington et al., 2014; Aumüller et al., 2020) 100 1.1M Word Embeddings GIST1M (Jegou et al., 2010b) 960 1M Spatial Image Descriptors
@@
Dataset $M$ $efC$ $efS$ Index Time Avg. Recall SIFT100M 32 500 500 23h 0.995 DEEP100M 32 500 750 20h 0.997 T2I100M 80 1000 2500 40h 0.970 GLOVE1M 16 500 500 2h 0.992 GIST1M 32 500 1000 6h 0.994
@@
Comparison Algorithms. We compare the results of DARTH with the Baseline we presented in Section 3.2.2. We also compare the performance of our approach against REM. The recall to efSearch mapping procedure is performed using 1K validation queries sampled from the learning sets of our datasets. Lastly, we compare our approach with the HNSW Learned Adaptive Early Termination (LAET) approach (Li et al., 2020). Note that LAET does not natively support declarative target recall with recall targets, since it is designed to terminate when all the nearest neighbors of a query have been found. For each query, after a fixed amount of HNSW search, LAET predicts the total number of distance calculations needed for this query to find all nearest neighbors. This value is then multiplied by a (hand-tuned) hyperparameter (called $multiplier$) to ensure that the number of distance calculations is sufficient. This hyperparameter tuning is performed using 1K validation queries sampled from the learning sets of our datasets. Then, the HNSW search terminates after the indicated distance calculations are performed. To achieve declarative recall with LAET, we manually tune the $multiplier$ to adjust the performance for each desired target recall $R_{t}$. Note that this implementation is not discussed in the original paper. During query answering, all algorithms use only a single core to answer each query, but multiple queries are executed in parallel, exploiting all available cores.
@@ 4.1. Training and Tuning
We analyzed the importance scores of the features used across all our datasets and values of $k$ (on average). The importance score expressed as a percentage of the total feature importance, was extracted from our GBDT recall predictor. Our analysis revealed that the features with the highest importance scores are $nstep$, $closestNN$, $firstNN$, $ninserts$, and $var$ (with importance scores of 16%, 16%, 16%, 14%, and 12%, respectively). This highlights that the estimation of the current recall is influenced by various search features, including the extent of the graph explored in the HNSW search, the nearest neighbors identified so far, and the initial nearest neighbor found at the beginning of the search.
@@
We conducted a feature ablation study to evaluate the performance of our recall predictor when using different combinations of input feature types from Table 1. Specifically, we compared the average validation $MSE$, $MAE$, and $R^{2}$ across all values of $k$ for various feature combinations for our datasets. The results indicate that using only the Index Metrics features yields moderate performance, with an $MSE$ of 0.0043, $MAE$ of 0.0318, and $R^{2}$ of 0.83. Incorporating either NN Distances or NN Stats alongside the Index Metrics improves the predictor’s performance, both achieving an $MSE$ of 0.0030, $MAE$ around 0.0269–0.0275, and $R^{2}$ of 0.88. In contrast, using NN Distances and NN Stats without Index Metrics leads to significantly worse results, with $MSE$ values exceeding 0.0191 and $R^{2}$ dropping below 0.30. As anticipated from the feature importance analysis, the most effective feature combinations involve both Index Metrics and at least one of the NN-based features. The overall best performance is achieved when all available features are used together, resulting in an $MSE$=0.0030, $MAE$=0.0269, and $R^{2}$=0.88. Consequently, our final recall predictor leverages the complete set of input features.
@@ 4.1.5. Recall Predictor Model Selection
For the grid-search, we report the results of two methods: Adaptive prediction interval tuning and a Static approach (i.e., with a fixed prediction interval, $mpi=ipi$). These methods are labeled as Adaptive-Grid-Search and Static-Grid-Search, respectively, and in our legends we refer to them as Ad-GS and St-GS for brevity. In each experiment, we selected the $mpi$ and $ipi$ configurations that achieved the best search times. We compared the grid-search methods to our heuristic hyperparameter selection method, described in Section 3.2.2, which is labeled Adaptive-Heuristic, and as Ad-Heur in our legends. To provide a comprehensive ablation study of the hyperparameter selection method, we also present results from a variant of the heuristic-based approach that does not employ adaptive prediction intervals, using fixed values of $ipi=mpi=\frac{dists_{R_{t}}}{4}$ (we selected to divide by 4 because this result gave us the best performance for this variant). We label this variant as Adaptive-Static, and in our legends we present it as Ad-St. Figure 6 illustrates the speedup achieved by each hyperparameter selection method across all datasets, for $R_{t}=0.90$ (Figure 4(a)) and $R_{t}=0.99$ (Figure 4(b)), using $k=50$. Both graphs show that the Adaptive methods outperform the corresponding Static methods, being up to 10% faster for the grid-search and up to 13% faster for the Heuristic method, while the Adaptive-Grid-Search method is the best-performing across all configurations. This is attributed to the adaptivity of the prediction intervals combined with the extensive hyperparameter tuning, resulting in excellent search times. Nevertheless, our Adaptive-Heuristic method, which does not involve any tuning at all, delivers comparable execution times (Adaptive-Grid-Search is only 5% faster). In DARTH, we automatically set the hyperparameter values using the Adaptive-Heuristic method, thus avoiding tuning all-together.
@@
Figure 3. $MSE$ for a varying number of training queries.
@@ 4.2.1. Recall Predictor Performance
We begin by presenting our recall predictor’s performance across the default testing query workloads of our datasets. The $MSE$, $MAE$, and $R^{2}$ measures are averaged over all $k$ values (we average to present the overall performance across all configurations), and are calculated by invoking the recall predictor at every point of the search for each query to examine the quality of the predictions fairly. The results are summarized in Table 5. The findings indicate that for all datasets, our models achieve very low $MSE$ and $MAE$ values, while maintaining high $R^{2}$ scores, demonstrating their effectiveness in estimating the recall of individual queries at any search stage.
Dataset $MSE$ $MAE$ $R^{2}$ SIFT100M 0.0029 0.0285 0.90 DEEP100M 0.0028 0.0270 0.88 GLOVE1M 0.0027 0.0189 0.87 GIST1M 0.0031 0.0307 0.88
@@ 4.2.5. Competitor Tuning Overheads
We now proceed to compare DARTH with competitor approaches. We note that DARTH is the only approach that natively supports declarative recall through early termination for any recall target $R_{t}$. In addition, REM also natively supports declarative recall for any recall target through the recall to efSearch mapping procedure it encapsulates. In contrast, LAET (with a tuned $multiplier$), the only related approach that uses early termination, requires specific tuning for each distinct $R_{t}$. Consequently, comparing LAET with DARTH necessitated extensive tuning for each recall target.
To fine-tune LAET for each $R_{t}$, we first performed a random search to identify the applicable ranges for the $multiplier$. We then employed binary search (due to the monotonic nature of the functions involved) to fine-tune the parameters. Specifically, we searched for $multiplier\in{0.10,0.15,0.20,\dots,3.00}$ and we evaluated the average recall values using a validation query set of 1K queries (same as the validation set of DARTH). The ranges and step sizes for the $multiplier$ were determined based on the results of the initial random search, which established the lower and upper bounds for the hyperparameter values of LAET. This limitation of the existing early termination method of LAET to address the problem of declarative recall highlights an important advantage of DARTH, which can directly start answering queries without the need for tuning. Figure 9 reports how many queries DARTH can answer before LAET finishes their tuning for $k=50$, demonstrating that DARTH is able to answer thousands of queries before LAET is tuned. Specifically, our approach can answer on average 6K, and up to 10K queries before LAET is tuned.
@@ 4.2.8. Comparison of DARTH with HNSW/REM Tuned for Hard Workloads
We now study the performance of DARTH for the T2I100M dataset, which contains OOD queries. We follow the same procedure as the other datasets, generating training data from 10K training queries originating from the learning set provided with the dataset. The vectors of the learning set follow the same distribution as the index (dataset) vectors. The training data generation time was 55 minutes, resulting in 340M training samples. Due to the bigger dataset search parameters, we logged a training sample every 2 distance calculations (instead of 1, like the rest of the datasets) to make sure that our training dataset size has a manageable size. The training time of the recall predictor was 320 seconds, and it achieved $MSE$=0.029, $MAE$=0.079, and $R^{2}$=0.54, by testing the predictor on 1K OOD queries from the default workload of the dataset. As expected, these results are not as good as those for the rest of the datasets (due to the multimodal nature of T2I100M), yet, they demonstrate the ability of the DARTH recall predictors to achieve good accuracy for OOD query workloads, just like they do for noisy workloads.
@@ 4.2.10. Extensions to IVF
After creating the plain IVF index, we executed 10K training queries to generate the training data for our IVF recall predictor. Note that, since IVF performs many more distance calculations for each query compared to HNSW, we had to reduce the logging frequency of our training data, gathering a training sample every 20 distance calculations for GLOVE1M and GIST1M, and every 50 distance calculations for DEEP100M and SIFT100M. This resulted in 315M training samples for SIFT100M, 310M for DEEP100M, 100M for GLOVE1M, and 133M for GIST1M. We trained a GBDT recall predictor, which achieved an average $MSE$=0.003 across all datasets, for the 1K testing queries of the default workloads.
@@ Acknowledgments
Supported by EU Horizon projects AI4Europe ($101070000$), TwinODIS ($101160009$), ARMADA ($101168951$), DataGEMS ($101188416$), RECITALS ($101168490$), and by $Y\Pi AI\Theta A$ NextGenerationEU project HARSH ($Y\Pi 3TA-0560901$).

### v2

@@ 3.3.2. Integration in IVF
We discuss the implementation of DARTH for the IVF (Douze et al., 2024) index as well, a popular Tree-based ANNS index. IVF performs k-means clustering over the vector collection, generating nlistnlist centroids. Each centroid operates as a bucket, and the collection vectors are placed in the bucket of their nearest centroid. IVF searches through the vectors of the nearest nprobenprobe cluster buckets to search for the nearest neighbors of a query vector.
DARTH can be effectively used for IVF with minimal changes to the input features of Table 1. Specifically, in DARTH for IVF, the $firstNN$ input feature represents the distance of the query to the closest centroid, while the nstepnstep feature represents the number of the cluster bucket we are currently searching. All other input features, as well as the dynamic recall predictor invocations with adaptive intervals, are the same as the HNSW implementation.
@@ 4. Experimental Evaluation
Setup. We conduct our experimental evaluation on a server with Intel® Xeon® E5-2643 v4 CPUs @ 3.40GHz (12 cores, 24 hyperthreads) and 500GB of available main memory. All algorithms are implemented in C/C++, embedded in the FAISS (Douze et al., 2024) library, with SIMD [^2] Single Instruction Multiple Data (SIMD): A parallel computing method where a single instruction operates simultaneously on multiple data points. support for the Euclidean Distance calculations. Our predictor models are implemented using the LightGBM (Ke et al., 2017) library. All implementations are compiled using g++ 11.4.0 on Ubuntu 22.04.4. We make our code publicly available on GitHub (dar, 2025).
@@
Dataset Dimension Base Vectors Description SIFT100M (Jégou et al., 2011) 128 100M Image Descriptors DEEP100M (Babenko and Lempitsky, 2016) 96 100M Deep Image Embeddings T2I100M (Simhadri et al., 2022) 200 100M Image and Text Embeddings GLOVE1M (Pennington et al., 2014; Aumüller et al., 2020) 100 1.1M Word Embeddings GIST1M (Jégou et al., 2011) 960 1M Spatial Image Descriptors
@@
Dataset $M$ efCefC efSefS Index Time Avg. Recall SIFT100M 32 500 500 23h 0.995 DEEP100M 32 500 750 20h 0.997 T2I100M 80 1000 2500 40h 0.970 GLOVE1M 16 500 500 2h 0.992 GIST1M 32 500 1000 6h 0.994
@@
Comparison Algorithms. We compare the results of DARTH with the Baseline we presented in Section 3.2.2. We also compare the performance of our approach against REM. The recall to efSearch mapping procedure is performed using 1K validation queries sampled from the learning sets of our datasets. Lastly, we compare our approach with the HNSW Learned Adaptive Early Termination (LAET) approach (Li et al., 2020). Note that LAET does not natively support declarative target recall with recall targets, since it is designed to terminate when all the nearest neighbors of a query have been found. For each query, after a fixed amount of HNSW search, LAET predicts the total number of distance calculations needed for this query to find all nearest neighbors. This value is then multiplied by a (hand-tuned) hyperparameter (called multipliermultiplier) to ensure that the number of distance calculations is sufficient. This hyperparameter tuning is performed using 1K validation queries sampled from the learning sets of our datasets. Then, the HNSW search terminates after the indicated distance calculations are performed. To achieve declarative recall with LAET, we manually tune the multipliermultiplier to adjust the performance for each desired target recall $R_{t}$. Note that this implementation is not discussed in the original paper. During query answering, all algorithms use only a single core to answer each query, but multiple queries can be executed in parallel, exploiting all available cores.
@@ 4.1. Training and Tuning
We analyzed the importance scores of the features used across all our datasets and values of $k$ (on average). The importance score expressed as a percentage of the total feature importance, was extracted from our GBDT recall predictor. Our analysis revealed that the features with the highest importance scores are nstepnstep, closestNNclosestNN, firstNNfirstNN, ninsertsninserts, and varvar (with importance scores of 16%, 16%, 16%, 14%, and 12%, respectively). This highlights that the estimation of the current recall is influenced by various search features, including the extent of the graph explored in the HNSW search, the nearest neighbors identified so far, and the initial nearest neighbor found at the beginning of the search.
@@
We conducted a feature ablation study to evaluate the performance of our recall predictor when using different combinations of input feature types from Table 1. Specifically, we compared the average validation $MSE$, MAEMAE, and $R^{2}$ across all values of $k$ for various feature combinations for our datasets. The results indicate that using only the Index Metrics features yields moderate performance, with an $MSE$ of 0.0043, $MAE$ of 0.0318, and $R^{2}$ of 0.83. Incorporating either NN Distances or NN Stats alongside the Index Metrics improves the predictor’s performance, both achieving an $MSE$ of 0.0030, $MAE$ around 0.0269–0.0275, and $R^{2}$ of 0.88. In contrast, using NN Distances and NN Stats without Index Metrics leads to significantly worse results, with $MSE$ values exceeding 0.0191 and $R^{2}$ dropping below 0.30. As anticipated from the feature importance analysis, the most effective feature combinations involve both Index Metrics and at least one of the NN-based features. The overall best performance is achieved when all available features are used together, resulting in an MSEMSE=0.0030, MAEMAE=0.0269, and $R^{2}$=0.88. Consequently, our final recall predictor leverages the complete set of input features.
@@ 4.1.5. Recall Predictor Model Selection
For the grid-search, we report the results of two methods: Adaptive prediction interval tuning and a Static approach (i.e., with a fixed prediction interval, $mpi=ipi$). These methods are labeled as Adaptive-Grid-Search and Static-Grid-Search, respectively, and in our legends we refer to them as Ad-GS and St-GS for brevity. In each experiment, we selected the mpimpi and ipiipi configurations that achieved the best search times. We compared the grid-search methods to our heuristic hyperparameter selection method, described in Section 3.2.2, which is labeled Adaptive-Heuristic, and as Ad-Heur in our legends. To provide a comprehensive ablation study of the hyperparameter selection method, we also present results from a variant of the heuristic-based approach that does not employ adaptive prediction intervals, using fixed values of $ipi=mpi=\frac{dists_{R_{t}}}{4}$ (we selected to divide by 4 because this result gave us the best performance for this variant). We label this variant as Adaptive-Static, and in our legends we present it as Ad-St. Figure 6 illustrates the speedup achieved by each hyperparameter selection method across all datasets, for $R_{t}=0.90$ (Figure 4(a)) and $R_{t}=0.99$ (Figure 4(b)), using $k=50$. Both graphs show that the Adaptive methods outperform the corresponding Static methods, being up to 10% faster for the grid-search and up to 13% faster for the Heuristic method, while the Adaptive-Grid-Search method is the best-performing across all configurations. This is attributed to the adaptivity of the prediction intervals combined with the extensive hyperparameter tuning, resulting in excellent search times. Nevertheless, our Adaptive-Heuristic method, which does not involve any tuning at all, delivers comparable execution times (Adaptive-Grid-Search is only 5% faster). In DARTH, we automatically set the hyperparameter values using the Adaptive-Heuristic method, thus avoiding tuning all-together.
@@
Figure 3. MSEMSE for a varying number of training queries.
@@ 4.2.1. Recall Predictor Performance
We begin by presenting our recall predictor’s performance across the default testing query workloads of our datasets. The $MSE$, MAEMAE, and $R^{2}$ measures are averaged over all $k$ values (we average to present the overall performance across all configurations), and are calculated by invoking the recall predictor at every point of the search for each query to examine the quality of the predictions fairly. The results are summarized in Table 5. The findings indicate that for all datasets, our models achieve very low $MSE$ and $MAE$ values, while maintaining high $R^{2}$ scores, demonstrating their effectiveness in estimating the recall of individual queries at any search stage.
Dataset MSEMSE MAEMAE $R^{2}$ SIFT100M 0.0029 0.0285 0.90 DEEP100M 0.0028 0.0270 0.88 GLOVE1M 0.0027 0.0189 0.87 GIST1M 0.0031 0.0307 0.88
@@ 4.2.5. Competitor Tuning Overheads
We now proceed to compare DARTH with competitor approaches. We note that DARTH is the only approach that natively supports declarative recall through early termination for any recall target $R_{t}$. In addition, REM also natively supports declarative recall for any recall target through the recall to efSearch mapping procedure it encapsulates. In contrast, LAET (with a tuned multipliermultiplier), the only related approach that uses early termination, requires specific tuning for each distinct $R_{t}$. Consequently, comparing LAET with DARTH necessitated extensive tuning for each recall target.
To fine-tune LAET for each $R_{t}$, we first performed a random search to identify the applicable ranges for the multipliermultiplier. We then employed binary search (due to the monotonic nature of the functions involved) to fine-tune the parameters. Specifically, we searched for $multiplier\in{0.10,0.15,0.20,\dots,3.00}$ and we evaluated the average recall values using a validation query set of 1K queries (same as the validation set of DARTH). The ranges and step sizes for the multipliermultiplier were determined based on the results of the initial random search, which established the lower and upper bounds for the hyperparameter values of LAET. This limitation of the existing early termination method of LAET to address the problem of declarative recall highlights an important advantage of DARTH, which can directly start answering queries without the need for tuning. Figure 9 reports how many queries DARTH can answer before LAET finishes their tuning for $k=50$, demonstrating that DARTH is able to answer thousands of queries before LAET is tuned. Specifically, our approach can answer on average 6K, and up to 10K queries before LAET is tuned.
@@ 4.2.8. Comparison of DARTH with HNSW/REM Tuned for Hard Workloads
We now study the performance of DARTH for the T2I100M dataset, which contains OOD queries. We follow the same procedure as the other datasets, generating training data from 10K training queries originating from the learning set provided with the dataset. The vectors of the learning set follow the same distribution as the index (dataset) vectors. The training data generation time was 55 minutes, resulting in 340M training samples. Due to the bigger dataset search parameters, we logged a training sample every 2 distance calculations (instead of 1, like the rest of the datasets) to make sure that our training dataset size has a manageable size. The training time of the recall predictor was 320 seconds, and it achieved MSEMSE=0.029, MAEMAE=0.079, and $R^{2}$=0.54, by testing the predictor on 1K OOD queries from the default workload of the dataset. As expected, these results are not as good as those for the rest of the datasets (due to the multimodal nature of T2I100M), yet, they demonstrate the ability of the DARTH recall predictors to achieve good accuracy for OOD query workloads, just like they do for noisy workloads.
@@ 4.2.10. Extensions to IVF
After creating the plain IVF index, we executed 10K training queries to generate the training data for our IVF recall predictor. Note that, since IVF performs many more distance calculations for each query compared to HNSW, we had to reduce the logging frequency of our training data, gathering a training sample every 20 distance calculations for GLOVE1M and GIST1M, and every 50 distance calculations for DEEP100M and SIFT100M. This resulted in 315M training samples for SIFT100M, 310M for DEEP100M, 100M for GLOVE1M, and 133M for GIST1M. We trained a GBDT recall predictor, which achieved an average MSEMSE=0.003 across all datasets, for the 1K testing queries of the default workloads.
@@ Acknowledgments
Supported by EU Horizon projects TwinODIS ($101160009$), DataGEMS ($101188416$), and by $Y\Pi AI\Theta A$ & NextGenerationEU project HARSH ($Y\Pi 3TA-0560901$).